# E2-BKI: Evidential Ellipsoidal Bayesian Kernel Inference for Uncertainty-aware Gaussian Semantic Mapping

### v1

@@ VI CONCLUSIONS
We propose an uncertainty-aware semantic mapping framework that addresses the limitations of S-BKI through anisotropic Gaussian primitives. Our approach tackles semantic uncertainty through evidential deep learning, spatial uncertainty through ellipsoidal kernels, and observation uncertainty through local context aggregation and pruning. By representing local geometry and semantics through Gaussian primitives, our method enables uncertainty-aware semantic propagation of sparse and noisy observations. Extensive experiments demonstrate superior performance in mapping quality and uncertainty calibration. The framework’s robustness and versatility across different map representations highlight its practicality for autonomous systems. Future work includes extending to dynamic environments with temporal consistency modeling and developing image-guided Gaussian refinement for enhanced geometric fidelity.
@@
VII Appendix: Preliminaries
@@ VII-A Semantic Bayesian Kernel Inference
where $k$ is a kernel function that satisfies the following conditions:
@@
The variance $\mathrm{Var}[\hat{\theta}_{m}^{\psi_{m}}]$ serves as a proxy for the uncertainty of the estimate [9, 13].
As each observation is weighted by the kernel $k$, its design determines both the spatial influence and efficiency of updates. A common choice is to let $k$ decay with the distance $d=\|\mathbf{\hat{x}}_{m}-\mathbf{x}_{n}\|$ and truncate beyond a length scale $\ell$, thereby emphasizing nearby points while suppressing distant ones. In practice, the sparse kernel [21] is widely adopted in continuous mapping [8, 9, 12] (Fig. 8):
@@ VII-B Evidential Deep Learning
To incorporate predictive uncertainty into the mapping process, we adopt the Evidential Deep Learning (EDL) framework [22] for semantic segmentation. Given a 2D dataset $\mathcal{D}_{\text{2D}}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N_{\text{img}}}$ with images $\mathbf{x}_{i}\in\mathbb{R}^{H\times W\times 3}$ and pixel-wise labels $\mathbf{y}_{i}\in\{0,1\}^{H\times W\times C}$, EDL enables direct estimation of both class probabilities and their uncertainty. Importantly, the same formulation can also be applied to 3D semantic segmentation, where a dataset $\mathcal{D}_{\text{3D}}=\{\mathbf{x}_{l},\mathbf{y}_{l}\}_{l=1}^{N_{\text{LiDAR}}}$ consists of point clouds $\mathbf{x}_{l}\in\mathbb{R}^{L\times 4}$ with per-point semantic labels $\mathbf{y}_{l}\in\{0,1\}^{L\times C}$. In this work, we perform 2D semantic segmentation to leverage rich semantic information while considering annotation efficiency.
Unlike conventional classifiers that rely on a $\text{SoftMax}(\cdot)$ layer to output class probabilities, EDL predicts non-negative evidence $\mathbf{\tilde{e}}_{i}=[\tilde{e}_{i}^{1},\ldots,\tilde{e}_{i}^{C}]$ parameterizing a Dirichlet with $\tilde{\alpha}_{i}^{c}=\tilde{e}_{i}^{c}+1$, yielding both class probabilities and a predictive uncertainty via the belief mass $\mathbb{\tilde{M}}_{i}=(\,\{\tilde{b}^{c}_{i}\}^{C}_{c=1},\tilde{u}_{i}\,)$:
$\tilde{b}_{i}^{c}=\frac{\tilde{e}_{i}^{c}}{\tilde{S}_{i}},\ \tilde{u}_{i}=\frac{C}{\tilde{S}_{i}},\text{ where }\tilde{u}_{i}+\sum_{c=1}^{C}\tilde{b}_{i}^{c}=1.$ (16)
Here, $\tilde{b}_{i}^{c}$ is the belief for class $c$, $\tilde{S}_{i}=\sum^{C}_{c=1}\tilde{\alpha}_{i}^{c}$ denotes the total evidence strength aggregated across classes, and $\tilde{u}_{i}$ quantifies the overall predictive uncertainty. The expected class probability is
@@
$\mathcal{L}_{\text{reg}}(\mathbf{x}_{i},\mathbf{y}_{i})$ $=D_{\text{KL}}\bigl{[}\text{Dir}\left(\mathbf{\tilde{p}}_{i}|\bm{\hat{\alpha}}_{i}\right)||\text{Dir}\left(\mathbf{\tilde{p}}_{i}|\langle 1,\ldots,1\rangle\right)\bigr{]},$
@@
Once the evidential model is trained to minimize the loss in (20), the probability $\mathbf{\tilde{p}}_{i}$ and its uncertainty $\tilde{u}_{i}$ are reliably estimated for each input $\mathbf{x}_{i}$. Incorporating these estimates into the mapping process can enhance the performance of semantic mapping, even under perceptually challenging off-road environments.
@@ VII-C Evidential Semantic Mapping
Although S-BKI provides an effective framework for continuous semantic mapping, it assumes equal reliability across all input points, neglecting the varying uncertainty inherent in semantic predictions. To address this limitation, EBS [10] proposes an uncertainty-aware extension of S-BKI by adopting EDL to obtain predictive uncertainties. Specifically, EBS adopts probabilistic semantic representations and an adaptive kernel, resulting in more robust and consistent semantic fusion.
To incorporate uncertainty into the mapping process, EBS generalizes the input semantic points $\mathcal{D}=\{(\mathbf{x}_{n},\mathbf{y}_{n})\}^{N}_{n=1}$ into evidential points $\mathcal{D^{\prime}}=\{(\mathbf{x}_{n},\mathbf{p}_{n},u_{n})\}^{N}_{n=1}$, where each point consists of a 3D coordinate $\mathbf{x}_{n}$, a probability distribution $\mathbf{p}_{n}$ over the $C$ semantic classes, and a corresponding uncertainty estimate $u_{n}$. Leveraging the generalized probabilistic input $\mathbf{p}_{n}$, the likelihood at query point $\mathbf{\hat{x}}_{m}$ can naturally be formulated using the Continuous Categorical distribution [30]:
@@
To further leverage uncertainty estimates, EBS introduces an adaptive kernel that modulates the spatial influence of each observation according to its reliability. Building upon the sparse kernel in (15), this adaptive formulation takes the form:
@@
where $\beta$ and $\gamma$ are scaling hyperparameters. For parameter simplification, we set $\gamma=1$. This adaptive kernel discards highly uncertain observations exceeding a dynamic threshold $\mathrm{U}_{\text{thr}}$, which filters out the top $\tilde{u}$ percentile of predictions. The spatial influence of the remaining observations is then scaled according to their uncertainty, enabling more confident predictions to contribute over a broader area. This sequential strategy of filtering outliers and adaptively weighting contributions yields a more reliable semantic map.
VIII Appendix: Methods
VIII-A Gaussian Initialization
@@
Clustering Given the evidential points $\mathcal{D^{\prime}}=\{(\mathbf{x}_{n},\mathbf{p}_{n},u_{n})\}_{n=1}^{N}$ as input, our framework performs clustering to construct Gaussian primitives. For stability, the points are first partitioned according to their most likely semantic label,
$\mathcal{D}^{\prime}_{c}=\{(\mathbf{x}_{n},\mathbf{p}_{n},u_{n})\in\mathcal{D^{\prime}}\;|\;\arg\max_{c^{\prime}}p_{n}^{c^{\prime}}=c\},$ (25)
@@
so that points assigned to the same semantic class are grouped together. Within each partition $\mathcal{D}^{\prime}_{c}$, we apply K-Means++ clustering [24] to form spatially coherent subsets $\{\mathcal{C}_{j}\}_{j=1}^{J_{c}}$ where $J_{c}$ denotes the number of clusters for class $c$. We then pool the clusters across all classes into a unified set $\mathbb{C}=\{\mathcal{C}_{j}\}_{j=1}^{J}$ with $J=\sum_{c=1}^{C}J_{c}$, removing the class index for subsequent Gaussian construction.
@@
Aggregation Once the clusters are constructed, each cluster $\mathcal{C}_{j}$ is converted into a Gaussian primitive $\mathcal{G}_{j}=(\mathbf{g}_{j},\mathbf{p}_{j},u_{j})$, where the geometric component $\mathbf{g}_{j}=(\bm{\mu}_{j},\bm{\Sigma}_{j})$ models the local geometry through mean $\bm{\mu}_{j}$ and covariance $\bm{\Sigma}_{j}$. For efficient merging, we internally employ the Method of Moments (MoM) [31], which maintains the first moment $\mathbf{m}^{(1)}_{j}$, the second moment $\mathbf{M}^{(2)}_{j}$, and the normalization constant $\eta_{j}$ for each Gaussian. These quantities can be converted back to Gaussian parameters as:
$\bm{\mu}_{j}=\frac{1}{\eta_{j}}\mathbf{m}^{(1)}_{j},\quad\bm{\Sigma}_{j}=\frac{1}{\eta_{j}}\mathbf{M}^{(2)}_{j}-\bm{\mu}_{j}\bm{\mu}_{j}^{\top}.$ (26)
@@
Meanwhile, the semantic component $(\mathbf{p}_{j},u_{j})$ is obtained by aggregating the semantic predictions $(\mathbf{p}_{n},u_{n})$ from all points in $\mathcal{C}_{j}$ using the DST-based combination rules [25]. In this framework, each prediction is expressed as a belief mass $\mathbb{M}=\bigl{(}\{b^{c}\}_{c=1}^{C},\,u\bigr{)}$, where $b^{c}$ is the belief mass for class $c$ and $u$ represents the overall uncertainty. Given two belief masses $\mathbb{M}_{1}=\bigl{(}\{b^{c}_{1}\}_{c=1}^{C},\,u_{1}\bigr{)}$ and $\mathbb{M}_{2}=\bigl{(}\{b^{c}_{2}\}_{c=1}^{C},\,u_{2}\bigr{)}$, their combination $\mathbb{M}=\mathbb{M}_{1}\oplus\mathbb{M}_{2}$ is defined as:
$b^{c}=\frac{1}{1-\delta}(b^{c}_{1}b^{c}_{2}+b^{c}_{1}u_{2}+b^{c}_{2}u_{1}),\quad u=\frac{1}{1-\delta}u_{1}u_{2},$ (27)
@@
where $\delta=\sum_{x\neq y}b^{x}_{1}b^{y}_{2}$ quantifies the conflict between two belief masses. To apply the rule in (27), the evidential points are reformulated as belief masses according to (16) and (17).
VIII-B Gaussian Refinement
@@
Merging To improve geometric stability and computational efficiency, we merge locally consistent Gaussian primitives. For each Gaussian, all neighboring primitives within a radius $d_{L}$ are examined for semantic consistency. If they all belong to the same semantic class, those within the smaller radius $d_{S}$ are merged. By requiring semantic homogeneity in the larger neighborhood $d_{L}$ before merging within $d_{S}$, we ensure that refinement occurs only in reliably uniform regions. As shown in Table I, we set $d_{L}=5\ell$ and $d_{S}=\ell$, so that merging remains conservative and avoids excessive fusion.
When these conditions are satisfied, the geometric components $\mathbf{g}_{j}$ are updated using the MoM representation already introduced in (26), which enables efficient in-place merging by simply accumulating first and second moments together with the normalization constants [17]. The semantic components are fused using the DST-based combination rule in (27), ensuring robust integration of class probabilities and uncertainties.
Pruning To address semantic inconsistencies that mainly arise from uncertain distant observations, we adopt a relative pruning strategy. Inspired by the human visual system, which maintains initial impressions until contradicted by more reliable evidence, we evaluate the reliability of each Gaussian primitive $\mathcal{G}_{j}$ and prune it when a conflicting but more reliable primitive $\mathcal{G}_{i}$ exists, as defined in (6). The sensitivity of pruning is controlled by $\epsilon$, which can be tuned according to sensor or environmental characteristics; however, we fix $\epsilon=2.5$ across all experiments and find this setting to be robust across diverse conditions.
VIII-C Gaussian as an Ellipsoid
Extending the BKI framework to Gaussian primitives requires a distance metric between a primitive and a query point. A naive choice, such as using the Euclidean distance between the mean $\bm{\mu}_{j}$ and the query, fails to capture the anisotropic shape of the Gaussian. On the other hand, probabilistic measures like the Mahalanobis distance reflect the covariance structure but are less intuitive and require nontrivial parameter tuning. Therefore, we explicitly interpret each Gaussian primitive $\mathcal{G}_{j}$ as an anisotropic ellipsoid and define the distance as the shortest Euclidean projection from the query point to the ellipsoid boundary. Formally, the ellipsoid $\mathcal{E}_{j}$ is defined by a Mahalanobis distance threshold $\tau$ corresponding to a desired confidence level:

### v2

@@ VI CONCLUSIONS
We propose an uncertainty-aware semantic mapping framework that addresses the limitations of S-BKI through anisotropic Gaussian primitives. Our approach tackles semantic uncertainty through evidential deep learning, spatial uncertainty through ellipsoidal kernels, and observation uncertainty through local context aggregation and pruning. By representing local geometry and semantics through Gaussian primitives, our method enables uncertainty-aware semantic propagation of sparse and noisy observations. Extensive experiments demonstrate superior performance in mapping quality and uncertainty calibration. The framework’s robustness and versatility across different map representations highlight its practicality for autonomous systems. Future work includes improving uncertainty estimation to obtain more calibrated predictive uncertainties, refining the formal definitions of the uncertainties, and extending the uncertainty-aware mapping framework to open-set settings.
@@
VII Background: Evidential Semantic Mapping with Bayesian Kernel Inference
@@ VII-A Semantic Bayesian Kernel Inference
where $k(\cdot,\cdot)$ is a kernel function that satisfies the following conditions:
@@
The variance $\mathrm{Var}[\hat{\theta}_{m}^{\psi_{m}}]$ serves as a proxy for the uncertainty of the estimate [^1] For brevity and readability, $\psi_{m}$ is denoted as $\psi$ in the main text. [7, 34].
As each observation is weighted by the kernel $k(\cdot,\cdot)$, its design determines both the spatial influence and efficiency of updates. A common choice is to let $k(\cdot,\cdot)$ decay with the distance $d=\|\mathbf{\hat{x}}_{m}-\mathbf{x}_{n}\|$ and truncate beyond a length scale $\ell$, thereby emphasizing nearby points while suppressing distant ones. In practice, the sparse kernel [23] is widely adopted in continuous mapping [6, 7, 5] (Fig. 8):
@@ VII-B Evidential Deep Learning
To incorporate predictive uncertainty into the mapping process, we adopt the Evidential Deep Learning (EDL) framework [28] for semantic segmentation. Given a 2D dataset $\mathcal{D}_{\text{2D}}=\{(\mathbf{X}_{i},\mathbf{Y}_{i})\}_{i=1}^{N_{\text{2D}}}$ with images $\mathbf{X}_{i}\in\mathbb{R}^{H\times W\times 3}$ and pixel-wise labels $\mathbf{Y}_{i}\in\{0,1\}^{H\times W\times C}$, EDL enables direct estimation of both class probabilities and their uncertainty. Importantly, the same formulation can also be applied to 3D semantic segmentation, where a dataset $\mathcal{D}_{\text{3D}}=\{(\mathbf{X}_{l},\mathbf{Y}_{l})\}_{l=1}^{N_{\text{3D}}}$ consists of point clouds $\mathbf{X}_{l}\in\mathbb{R}^{L\times 4}$ with per-point semantic labels $\mathbf{Y}_{l}\in\{0,1\}^{L\times C}$. In this work, we perform 2D semantic segmentation to leverage rich semantic information while considering annotation efficiency. For notational convenience, we denote an arbitrary pixel and its corresponding label in $(\mathbf{X}_{i},\mathbf{Y}_{i})$ by $(\mathbf{x}_{i},\mathbf{y}_{i})$.
Unlike conventional classifiers that rely on a $\text{SoftMax}(\cdot)$ layer to output class probabilities, EDL predicts non-negative evidence $\mathbf{\tilde{e}}_{i}=[\tilde{e}_{i}^{1},\ldots,\tilde{e}_{i}^{C}]$ for each $\mathbf{x}_{i}$, parameterizing a Dirichlet $\text{Dir}(C,\tilde{\bm{\alpha}}_{i})$ with $\tilde{\alpha}_{i}^{c}=\tilde{e}_{i}^{c}+1$. EDL yields both class probabilities and a predictive uncertainty via the belief mass $\mathbb{\tilde{M}}_{i}=(\{\tilde{b}^{c}_{i}\}^{C}_{c=1},\tilde{u}_{i})$:
$\tilde{b}_{i}^{c}=\frac{\tilde{e}_{i}^{c}}{\tilde{S}_{i}},\quad\tilde{u}_{i}=\frac{C}{\tilde{S}_{i}},\quad\text{where }\tilde{u}_{i}+\sum_{c=1}^{C}\tilde{b}_{i}^{c}=1.$ (16)
Here, $\tilde{b}_{i}^{c}$ is the belief for the class $c$, $\tilde{S}_{i}=\sum^{C}_{c=1}\tilde{\alpha}_{i}^{c}$ denotes the total evidence strength aggregated across classes, and $\tilde{u}_{i}$ quantifies the overall predictive uncertainty. The expected class probability is
@@
$\mathcal{L}_{\text{reg}}(\mathbf{x}_{i},\mathbf{y}_{i})$ $=D_{\text{KL}}\Big[\text{Dir}\left(\mathbf{\tilde{p}}_{i}|\bm{\hat{\alpha}}_{i}\right)\,\|\,\text{Dir}\left(\mathbf{\tilde{p}}_{i}|[1,\ldots,1]\right)\Big],$
@@
Once the EDL model is trained to minimize the loss in (20), the probability $\mathbf{\tilde{p}}_{i}$ and its uncertainty $\tilde{u}_{i}$ are reliably estimated for each input $\mathbf{x}_{i}$. Incorporating these estimates into the mapping process can enhance the performance of semantic mapping, even under perceptually challenging off-road environments.
@@ VII-C Evidential Semantic Mapping
Although S-BKI provides an effective framework for continuous semantic mapping, it assumes equal reliability across all input points, neglecting the varying uncertainty inherent in semantic predictions. To address this limitation, we introduced EBS [17], an evidential semantic mapping framework that extends S-BKI by explicitly incorporating predictive uncertainty via Evidential Deep Learning. Specifically, EBS adopts probabilistic semantic representations and an adaptive kernel, resulting in more robust and consistent semantic fusion.
To incorporate uncertainty into the mapping process, EBS generalizes the input semantic points $\mathcal{D}=\{(\mathbf{x}_{n},\mathbf{y}_{n})\}^{N}_{n=1}$ into evidential points $\mathcal{D^{\prime}}=\{(\mathbf{x}_{n},\mathbf{p}_{n},u_{n})\}^{N}_{n=1}$, where each point consists of a 3D coordinate $\mathbf{x}_{n}$, a probability distribution $\mathbf{p}_{n}$ over the $C$ semantic classes, and a corresponding uncertainty estimate $u_{n}$. Leveraging the generalized probabilistic input $\mathbf{p}_{n}$, the likelihood at query point $\mathbf{\hat{x}}_{m}$ can be formulated naturally using the Continuous Categorical distribution [9]:
@@
To further leverage uncertainty estimates, EBS introduces an adaptive kernel that modulates the spatial influence of each observation according to its reliability. Building upon the sparse kernel in (15), the adaptive formulation takes the form:
@@
where $\beta$ and $\gamma$ are scaling hyperparameters [^2] To simplify the parameterization, we fix $\gamma=1$, and we set $\sigma_{0}=1$ to satisfy (11).. This adaptive kernel discards highly uncertain observations exceeding a dynamic threshold $\mathrm{U}_{\text{thr}}$, which filters out the top $\tilde{u}$ percentile of each input. The spatial influence of the remaining observations is then scaled according to their uncertainty, enabling more confident predictions to contribute over a broader region. This strategy of filtering outliers and adaptively weighting contributions yields a more reliable semantic map.
VIII Evidential Ellipsoidal BKI
VIII-A Gaussian Initialization (Section IV-B)
@@
Clustering Our framework performs clustering to construct Gaussian primitives from the evidential points $\mathcal{D^{\prime}}=\{(\mathbf{x}_{n},\mathbf{p}_{n},u_{n})\}_{n=1}^{N}$. For clustering stability, the points are first partitioned by their most likely semantic label:
$\mathcal{D}^{\prime}_{c}=\{(\mathbf{x}_{n},\mathbf{p}_{n},u_{n})\in\mathcal{D^{\prime}}\;|\;\arg\max_{c^{\prime}}p_{n}^{c^{\prime}}=c\}.$ (25)
@@
Clustering is thus performed independently within each semantic class, as we empirically found this strategy to yield more stable and reliable performance. Within each partition $\mathcal{D}^{\prime}_{c}$, we apply K-Means++ clustering [1] to form spatially coherent subsets $\{\mathcal{C}_{j}\}_{j=1}^{J_{c}}$ where $J_{c}$ denotes the number of clusters for class $c$. We then pool the clusters across all classes into a unified set $\mathbb{C}=\{\mathcal{C}_{j}\}_{j=1}^{J}$ with $J=\sum_{c=1}^{C}J_{c}$, removing the class index for subsequent Gaussian construction.
@@
Aggregation Each cluster $\mathcal{C}_{j}$ is then converted into a Gaussian primitive $\mathcal{G}_{j}=(\bm{\mu}_{j},\bm{\Sigma}_{j},\mathbf{p}_{j},u_{j})$, where the geometric component $(\bm{\mu}_{j},\bm{\Sigma}_{j})$ models the local geometry through mean $\bm{\mu}_{j}$ and covariance $\bm{\Sigma}_{j}$, and the semantic component $(\mathbf{p}_{j},u_{j})$ represents the evidential beliefs.
For efficient merging, we employ the Method of Moments (MoM) [26] to represent each Gaussian. MoM represents each Gaussian using the unnormalized first moment $\mathbf{m}^{(1)}_{j}$, the unnormalized second moment $\mathbf{M}^{(2)}_{j}$, and a normalization constant $w_{j}$. These quantities can be converted back to Gaussian parameters as:
$\bm{\mu}_{j}=\frac{1}{w_{j}}\mathbf{m}^{(1)}_{j},\quad\bm{\Sigma}_{j}=\frac{1}{w_{j}}\mathbf{M}^{(2)}_{j}-\bm{\mu}_{j}\bm{\mu}_{j}^{\top}.$ (26)
Using this representation, Gaussian primitives can be efficiently merged by directly accumulating their moments. Specifically, when two primitives $\mathcal{G}_{A}$ and $\mathcal{G}_{B}$ are merged to yield $\mathcal{G}$, the merged normalization constant, the first moment, and the second moment are obtained as:
$\begin{gathered}\mathbf{m}^{(1)}=\mathbf{m}^{(1)}_{A}+\mathbf{m}^{(1)}_{B},\quad\mathbf{M}^{(2)}=\mathbf{M}^{(2)}_{A}+\mathbf{M}^{(2)}_{B},\\ w=w_{A}+w_{B}.\end{gathered}$ (27)
The resulting Gaussian parameters $\bm{\mu}$ and $\bm{\Sigma}$ are then recovered using (26). This moment-based merging is computationally efficient and well-suited for incremental map construction.
@@
Meanwhile, the semantic component $(\mathbf{p}_{j},u_{j})$ is obtained by aggregating the semantic predictions $(\mathbf{p}_{n},u_{n})$ from all points in $\mathcal{C}_{j}$ using the DST-based combination rules [11]. In this framework, each prediction is expressed as a belief mass $\mathbb{M}=\bigl(\{b^{c}\}_{c=1}^{C},\,u\bigr)$, where $b^{c}$ is the belief mass for class $c$ and $u$ represents the overall uncertainty. Given two belief masses $\mathbb{M}_{A}=\bigl(\{b^{c}_{A}\}_{c=1}^{C},\,u_{A}\bigr)$ and $\mathbb{M}_{B}=\bigl(\{b^{c}_{B}\}_{c=1}^{C},\,u_{B}\bigr)$, their combination $\mathbb{M}=\mathbb{M}_{A}\oplus\mathbb{M}_{B}$ is defined as:
$b^{c}=\frac{1}{1-\eta}(b^{c}_{A}b^{c}_{B}+b^{c}_{A}u_{B}+b^{c}_{B}u_{A}),\quad u=\frac{1}{1-\eta}u_{A}u_{B},$ (28)
@@
where $\eta=\sum_{x\neq y}b^{x}_{A}b^{y}_{B}$ quantifies the conflict between two belief masses. To apply the rule in (28), the evidential points are converted into belief masses using (16) and (17).
VIII-B Gaussian Refinement (Section IV-C)
@@
Merging To improve geometric stability and computational efficiency, we merge locally consistent Gaussian primitives. For each Gaussian, all neighboring primitives within a radius $d_{L}$ are examined for semantic consistency. If they all belong to the same semantic class, those within the smaller radius $d_{S}$ are merged. By enforcing semantic homogeneity over the larger neighborhood $d_{L}$ prior to merging within $d_{S}$, we ensure that refinement occurs only in reliably uniform regions. As summarized in Table I, we set $d_{L}=5\ell$ and $d_{S}=\ell$, which yields a conservative merging strategy and prevents excessive fusion.
When these conditions are satisfied, the geometric components $(\bm{\mu}_{j},\bm{\Sigma}_{j})$ are updated using the MoM representation introduced in (26) and (27). This formulation enables efficient in-place merging through moment accumulation [20]. The semantic components are fused using the DST-based combination rule in (28), ensuring robust integration of class probabilities and uncertainties.
Pruning To address semantic inconsistencies that mainly arise from uncertain distant observations (as shown in Fig. 3), we adopt a relative pruning strategy. Inspired by the human visual system, which maintains initial impressions until contradicted by more reliable evidence, we evaluate the reliability of each Gaussian primitive $\mathcal{G}_{j}$ and prune it when a conflicting but more reliable primitive $\mathcal{G}_{i}$ exists, as defined in (6). The sensitivity of pruning is controlled by $\epsilon$, which can be tuned according to sensor or environmental characteristics; however, we fix $\epsilon=2.5$ across all experiments and find this setting to be robust across diverse conditions.
VIII-C Gaussian as an Ellipsoid (Section IV-D)
When BKI inference is formulated over Gaussian primitives rather than raw data points, a new notion of distance between a primitive and a query location becomes necessary. In conventional BKI, kernel inference operates on individual observations using Euclidean distance, whereas our framework introduces Gaussian primitives as an intermediate scene representation. Consequently, kernel inference must be defined between a Gaussian primitive and a query point.
An appropriate distance metric must capture not only the location $\bm{\mu}_{j}$ of a Gaussian primitive $\mathcal{G}_{j}$ but also its spatial extent and anisotropic geometry $\bm{\Sigma}_{j}$. A naive choice, such as the Euclidean distance between the primitive mean $\bm{\mu}_{j}$ and the query point, fails to capture the anisotropic shape of the Gaussian. Probabilistic measures such as the Mahalanobis distance incorporate the covariance information but are less intuitive and require nontrivial parameter tuning. Therefore, we explicitly interpret each Gaussian primitive $\mathcal{G}_{j}$ as an anisotropic ellipsoid and define the distance as the shortest Euclidean projection from the query point to the ellipsoid boundary. Formally, the ellipsoid $\mathcal{E}_{j}$ is defined by a Mahalanobis distance threshold $\tau$ corresponding to a desired confidence level: